Shift Summary for Nov. 28 Day shift

Sydney A. Taegar (taegar@kaon1.physics.arizona.edu)
Sun, 28 Nov 1999 14:23:29 -0700 (MST)

Shift Summary for Day shift, 11/28/99

On Shift: S. Taegar, U. Erkarsalan
+ cast of tens

Hours of beam delivered at all intensities: 0.8
Hours of beam delivered at nominal intensity: 0.8
Hours of 99 data (special runs not included): 0.8
Hours of special runs (laser scans, muon runs...): 0

=========================================================================
Runs ended this shift:
15075 BEAM_799 2.1M events

=========================================================================
07:25 TRD HV alarm claims TRD HV changed by more than 10V.
Valeri says this is due to a HV readback error.

07:50 Power outage to fixed target and CHL (Central Helium Liquifier)
Jason H. and Ugur page experts and start powering systems
off. A tardy Sydney arrives shortly threafter.

09:40 Power returns.

09:50 Shift crew + various experts start bringing systems back up.

09:55 Power outage.

10:00 Power returns. Gun-shy ktev'ers decide to wait for
situation to stabilize before attempting recovery.

11:20 Vacuum system back to normal. Gate valve is closed -- needs
to be opened before securing KTeV hall.

11:40 Bob Mau calls to inform us that they have lost CHL (which
had been on backup power) and anticpate beam will not
return for at least 12 hrs and possibly as long as 24 hrs.

14:12 Feeder repair and reconfiguration is complete. CHL and
TeV recovery underway. Current estimate for HEP operation
is 165 hrs. Next update at 18:00.

Various experts paged, ktev recovery underway.

14:40 Doug Jensen called regarding vacuum gate valve. He is
due in at midnight and will reopen it then.

15:00 Recovery continues....